OpenStudy (anonymous):

9 An aqueous solution was prepared containing 1.0mol of AgNO3 and 1.0mol of FeSO4 in 1.00 dm 3 of water. When equilibrium was established, there was 0.44mol of Ag + (aq) in the mixture. Ag + (aq) + Fe 2+ (aq) Ag(s) + Fe 3+ (aq) What is the numerical value of Kc? A 0.35 B 0.62 C 1.62 D 2.89

OpenStudy (mos1635):

1 mole of AgNO3 gives 1 mole of Ag+ 1 mole of FeSO4 gives 1 mole of Fe2+ Ag + (aq) + Fe 2+ (aq) --> Ag(s) + Fe 3+ (aq) inital 1 1 reacts/rpoduce -x -x +x +x in equilbrium 1-x 1-x x x 1-x=0.44 x=0.56 Kc=[ Fe 3+ (aq) ]/[ Ag + (aq) ]*[ Fe 2+ (aq) ] ...... Kc=x*V/(1-x)^2 Kc=0.56*1/0.44^2 Kc=2.89

OpenStudy (anonymous):

thx :)